OpenStudy (howard-wolowitz):

Simple, but hard math question:

OpenStudy (howard-wolowitz):

Click to enlarge!

OpenStudy (anonymous):

i will help you

OpenStudy (howard-wolowitz):

well thank you

OpenStudy (anonymous):

your welcome :)

OpenStudy (solomonzelman):

divide every term by \(\sin^2\theta\)

OpenStudy (solomonzelman):

and then tell me what you get

OpenStudy (howard-wolowitz):

ok let me do that

OpenStudy (solomonzelman):

sure, take your time

OpenStudy (howard-wolowitz):

ok i got 1=1

OpenStudy (solomonzelman):

???

OpenStudy (howard-wolowitz):

if i did that correctly

OpenStudy (howard-wolowitz):

ok let me try again

OpenStudy (solomonzelman):

if you got 1=1 you did it correctly, but you went too far

OpenStudy (solomonzelman):

I will direct you

OpenStudy (howard-wolowitz):

oh

OpenStudy (solomonzelman):

Your identity is: \(\large\color{slate}{\displaystyle \sin^2\theta+\cos^2\theta=1}\) and you are given some toher options, and you have to choose one option equivalent to that. I advised you to divide everything by \(\large\color{slate}{\displaystyle \sin^2\theta}\), and this is what I mean: \(\large\color{red}{\displaystyle \frac{\color{black}{\sin^2\theta}}{\sin^2\theta}\color{black}{+}\frac{\color{black}{\cos^2\theta}}{\sin^2\theta}\color{black}{=}\frac{\color{black}{1}}{\sin^2\theta}}\)

OpenStudy (solomonzelman):

and then simplify this expression, and tell me what is the correct option.

OpenStudy (howard-wolowitz):

alright I got \[\csc ^{2} 0\]

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \large\color{red}{\displaystyle \frac{\color{black}{\sin^2\theta}}{\sin^2\theta}\color{black}{+}\frac{\color{black}{\cos^2\theta}}{\sin^2\theta}\color{black}{=}\frac{\color{black}{1}}{\sin^2\theta}}}\) \(\large\color{slate}{\displaystyle ~~~~\Downarrow ~~+~~~\Downarrow~~~=~~\Downarrow}\) \(\large\color{slate}{\displaystyle ~~~~~1~~+~~\cot^2\theta=\csc^2\theta}\)

OpenStudy (howard-wolowitz):

i see. from ^ i choose B. i didnt realize that you meant solve the whole thing though

OpenStudy (howard-wolowitz):

my bad

OpenStudy (solomonzelman):

it's ol

OpenStudy (solomonzelman):

ok